OpenStudy (anonymous):

Given the functions f(x) = 3x - 1 and g(x) = 3x + 4, which operation results in the smallest coefficient on the x term?

OpenStudy (anonymous):

What about this questio too:Given the functions j(x) = x2 - 9 and k(x) = -x + 7, which operation results in a 3rd degree polynomial? help?!